OpenStudy (anonymous):

Anthony purchased a bedroom set for $3,860 using a six-month deferred payment plan with an interest rate of 27.29%. What is the balance after the deferment period if payments of $172 are made each month? a) $2,828.00 b) $4,417.57 c) $3,385.57 d) $3,860.00

jimthompson5910 (jim_thompson5910):

I'm not exactly sure how the deferred payment plan is set up, but I'm assuming that you don't make any payments for those 6 months and interest is being accrued (ie added on) Also, I'm assuming it's being compounded monthly If that's the case, then... A = P(1 + r/n)^(nt) in this case, P = 3860, r = 0.2729, n = 12 and t = 6/12 (t is in years) A = 3860*(1+0.2729/12)^(12*6/12) A = ??? Tell me what you get

OpenStudy (anonymous):

I got 4417.57

jimthompson5910 (jim_thompson5910):

me too, congrats

jimthompson5910 (jim_thompson5910):

So this is the amount that she must pay off after the 6 month deferred period is up (ie this is the balance after 6 months)

jimthompson5910 (jim_thompson5910):

where no payments were made during that 6 month deferred period

OpenStudy (anonymous):

Yep! Thank you for your help

jimthompson5910 (jim_thompson5910):

yw